# Distribution System Reconfiguration to Mitigate Load Altering Attacks via Stackelberg Games

### v5

@@ Abstract
The integration of IoT-controllable devices in power systems (such as smart electric vehicle charging stations, heat pumps, etc.), despite their benefits, raises novel cybersecurity concerns. Vulnerabilities in these devices can be leveraged to launch load-altering attacks (LAAs) that can potentially compromise the safety of power systems. In this paper, we analyze the impact of LAAs on the voltage profile of distribution networks (DNs). We first derive closed-form expressions to quantify the attacks’ impact. Using the insights derived from this analysis, we then propose a reactive defense method to mitigate LAAs based on reconfiguring the DNs. We also study optimal defense strategies that are robust to LAAs by exploiting non-cooperative sequential game theory. The proposed solution takes into account the potential uncertainties in the attack localization. Furthermore, we propose a Bayesian optimization (BO) approach to compute the Stackelberg equilibrium of the game, which reduces the computational burden. Our results show that attacks launched on the deepest nodes in the DN have the most detrimental effect on the grid voltage profile. Furthermore, the proposed game-theoretic strategy successfully mitigates the effect of the attack while ensuring minimum system reconfiguration.
Index Terms:
Distribution network, Cybersecurity, Load-altering attack (LAA), Reconfiguration, Stackelberg game, Bayesian optimization.
@@ I Introduction
Internet-of-Things (IoT) enabled devices offer enhanced end-user convenience, improved efficiency and flexibility to power systems for load peak management, which has driven a notable surge in their adoption. However, beyond their evident benefits, these devices also present potential vulnerabilities, serving as entry points for cyber attackers to exploit and compromise the security of power systems. Specifically, load-altering attacks (LAA) in power networks with high IoT-enabled device penetration pose a significant cybersecurity threat [2, 3, 4, 5].
The concept of LAAs was first introduced in [3] in which adversaries turn a group of IoT-controllable electrical loads into bots and turn them on/off simultaneously to harm the stability of the system. The manipulation of loads disrupts the balance between power generation and demand, leading to frequency instabilities in transmission networks [6, 4]. In distribution networks (DNs), LAA can result in elevated line flows, causing higher voltage drops leading to voltage constraint violation [7]. Furthermore, the data required to launch successful LAAs can be obtained from publicly available information [8] or circumvented [9].
@@ I-A Literature Survey
Attack Impact Analysis and Viability: Researchers in [4] investigated the impact of LAAs on transmission systems and identified several effects, including line failures, unsafe frequency deviations, disruption in grid restarting, and increased operational costs. The research in [10] examined the effects of LAA under a more realistic setting consisting of protection (such as $N-1$ security) and load-shedding schemes and showed that LAA can still cause outages and islands. [11] examined the feasibility and impact of LAAs under low demand and low inertia conditions experienced during the COVID-19 pandemic, and showed that under these conditions, it can be easier to launch LAAs that lead to unsafe grid conditions. A study on the impact of attacks on IoT-enabled devices on interconnected transmission and distribution systems was carried out in [12].
Reference [2] proposed the so-called dynamic LAA (DLAA), in which the adversary continuously toggles the compromised load devices on and off, guided by a feedback control loop in response to the system’s frequency fluctuations. In [5], an analytical framework was introduced to analyze the impact of LAA on transmission systems and identify the nodes from which an attacker can launch the most effective attacks using the theory of second-order dynamical systems. In [13], a rare-event sampling algorithm was proposed that uncovers the spatial and temporal distribution of impactful DLAAs while considering the security constraint of $N-1$.
The growing popularity of electric vehicles (EVs) and the spread of EV charging stations (EVCS) make them a potential target for the LAA. Reference [14] explores the weak points of EVCS, such as firmware flaws, vulnerabilities in management systems, and the security scarcity of mobile apps that allow adversaries to launch LAAs. Subsequently, to detect such attacks, they propose model-based and data-driven approaches. Furthermore, in [15] authors study the problem of launching DLAA on coordinated distributed EVCS, while considering potential communication delays when sending attack commands. Furthermore, their attack circumvents existing defensive solutions, such as traveling wave and transient energy-based methods. They employ a robust optimization method to launch a coordinated attack from multiple locations to maximize the impact of the attack.
Attack Mitigation: Another stream of research investigates the mitigation of LAAs. The existing mitigation methods can be categorized into: i) offline; or ii) online methods.
@@
Offline methods: Offline defenses try to install preventive measures to stop the impact of LAAs. For instance, [16] proposed algorithms to determine the operating points for generators in a way to prevent line overloads caused by potential botnet-type attacks against IoT devices. Reference [2] presented a mitigation framework based on securing the most critical loads. This method found the minimum magnitude of loads needed to be protected in order to guarantee frequency stability in the event of DLAAs. [21] proposed a zero-sum Stackelberg game formulation to install reactive power compensation (RPC) to reduce the impact of the attack. While the works above focus on transmission systems, [7] introduced a mitigation approach tailored specifically for DNs. Their research focuses on identifying optimal locations for deploying soft open points (SOPs) and refining their operation to mitigate the effects of attacks on voltage deviations.
@@
Online methods: Despite the effectiveness of the offline methods, these measures may be too costly as the preventive features must be enabled irrespective of whether an attack occurs or not (e.g., uneconomic generator operating points to cover for LAAs). Online methods, on the other hand, involve determining defensive actions to counter the effects of LAAs once the attack is launched, via reactive measures. In [17], a cyber-resilient economic dispatch method is introduced to mitigate LAAs based on altering the frequency droop control parameters of inverter-based resources to counter the destabilizing effects of LAAs. [18] proposed a framework in which electric vehicles are designed as feedback controllers that can mitigate the impact of LAA based on $H-2$ and $H-\infty$ norms. To analyze the manoeuvres of a strategic attacker initiating DLAA, [19] introduced a multi-stage game approach. In this game, the defensive actions involved load shedding, and the ultimate objective is to achieve a strategic balance between DLAA and the necessary amount of load shedding, reaching a Nash equilibrium (NE). [20] formulated the cybersecurity of the power system in both cyber and physical layers as a game-theoretic formulation.
Reconfiguration of DNs: Reconfiguration of DNs is a well-studied topic in which the topology of the networks is changed by turning existing line switches on/off [22]. The existence of switches in lines which also can be used to isolate potential faults in lines provides a paramount amount of possible configurations for the grid. This usually takes place for reducing power loss, network balancing, voltage profile rectification, enhancing network restoration, and network reliability [23, 24, 25]. The fact that the operator can perform the changes in topology swiftly, only by sending on/off commands to switches makes this capability a great fit for responding to a potential threat.
@@ I-B Research Gap and Contributions
Based on the discussions of Section I-A and Table I it can be inferred that there is a dearth of studies for analyzing the impact of LAA on DNs. Similarly, the majority of mitigations are designed to compensate for attacks in transmission grids. This paper addresses this gap in LAA studies in DNs. Furthermore, this paper is the first to consider mitigation against strategic attackers in the context of LAAs on DNs. Drawing from the insights presented in [26, 27], we also considered noisy system measurements and uncertainties associated with attack detection/localization in our mitigation scheme.
We start with a rigorous investigation of the impact of LAAs on DNs. We integrate the ZIP load model [28] into our formulations to obtain a realistic attack impact of DNs. Our work provides analytical insights into the impact of LAAs and identifies the most vulnerable nodes of the system. The results are then utilized to guide the design of the mitigation strategy. To take a strategic attacker into account and to defend against the worst-case attacks, we model the attacker-defender interaction using a Stackelberg game framework with the attacker as the leader and the defender as the follower. This sequential game is relevant for reactive defenses, in which mitigation is triggered only when an attack is detected. To obtain the players’ payoffs, the reconfiguration problem should be solved repeatedly. We reformulate this problem as a mixed integer linear programming (MILP) optimization by using the linearized distribution flow (LinDistFlow) and the ZP approximations. The resulting formulation provides a significant computational speed-up in calculating the game’s payoffs.
Despite simplifying the optimization, obtaining the Stackelberg equilibrium remains computationally complex in cases involving large sets of player actions since it requires repeatedly solving optimization problems to compute the payoff functions (e.g., if an attacker aims to target $n_{a}$ buses from a pool of $N$ candidate victim buses, the total number of possible attacks is given by $\binom{N}{n_{a}}$, which can be very large.). To mitigate this burden, we propose using Bayesian optimization (BO). Rather than exhaustively evaluating all possible actions, this approach constructs a probabilistic model and iteratively refines the model, focusing on the most promising actions to identify optimal strategies swiftly.
This paper significantly extends our preliminary work [1] in terms of using the insights derived from our attack impact analysis to design a mitigation strategy to counter LAAs. The key contributions are summarized as follows:
$\bullet$ Deriving closed-form expressions of the bus voltages of the DN in the presence of voltage-dependent loads.
$\bullet$ Analysing the effects of LAAs launched at different nodes of the DN and obtaining the minimum number of IoT-controllable devices required for a detrimental attack.
$\bullet$ Proposing a novel sequential game to counter the effect of LAAs using grid reconfiguration, considering the potential errors and uncertainties associated with localizing the attack.
$\bullet$ Using BO to obtain the game’s Stackelberg equilibrium to reduce its computational complexity.
@@ II-A DN Model
The DN under study is represented by the connected directed graph $G=\{\mathcal{N},\mathcal{L}\}$, where $\mathcal{N}=\{1,2,\ldots,N\}$ denotes the set of buses and $\mathcal{L}$ denotes the set of branches. This graph has a radial structure, hence it is a tree. Except for bus $1$, which is the root, each bus is referred to as the ‘child’ of its parent bus which is the adjacent bus closer to bus $1$ by one branch. Thus, the set of branches is defined as =$\mathcal{L}=\{(\pi_{i},i)\ |\ \pi_{i},\ i\in\mathcal{N}\}$, where $\pi_{i}$ represents the parent of bus $i$. In this configuration, bus $1$ represents the generator bus. We denote by $\mathcal{D}_{k}$ the set of buses which forms the unique path connecting bus $1$ to bus $k$, excluding bus $1$ and including bus $k$. The depth of each bus represents the distance in terms of the number of branches between that bus and the root bus.
@@ II-B1 ZIP Load Model
ZIP and exponential load models are the most implemented ones in the industry [29]. Experimental values for voltage dependency of loads are captured and fitted in the ZIP model in [30], which we integrate into our formulations. The power demand under the ZIP load model is given in [28] as follows:
$s^{zip}_{i}(v_{i})=p^{l}_{i}(\alpha_{p}+\beta_{p}v_{i}+\gamma_{p}v^{2}_{i})+jq^{l}_{i}(\alpha_{q}+\beta_{q}v_{i}+\gamma_{q}v^{2}_{i}),$ (1)
where $\alpha_{k}+\beta_{k}+\gamma_{k}=1$, where $k=\{p,q\}.$ Furthermore, $v_{i}$ is the nodal voltage while $p^{l}_{i}$ and $q^{l}_{i}$ are active and reactive load demands of bus $i$ at rated voltage. The ZIP load model captures the voltage dependency of real-world loads.
@@ II-B2 ZP Approximation
Based on (1), the ZIP model is a function of both $v_{i}$ and $v^{2}_{i}$. This causes the optimization tasks involving the power flow in the presence of ZIP loads to become nonconvex and complex. To tackle this problem, [31] have provided an approximate model for ZIP loads given by
$s^{zp}_{i}(v_{i})=p^{l}_{i}(\alpha^{\prime}_{p}+\gamma^{\prime}_{p}v_{i}^{2})+jq^{l}_{i}(\alpha^{\prime}_{q}+\gamma^{\prime}_{q}v^{2}_{i}),$ (2)
where $\alpha^{\prime}_{p}=\alpha_{p}+\frac{\beta_{p}}{2}$ ,$\alpha^{\prime}_{q}=\alpha_{q}+\frac{\beta_{q}}{2}$ ,$\gamma^{\prime}_{p}=\gamma_{p}+\frac{\beta_{p}}{2}$ , and $\alpha^{\prime}_{q}=\gamma_{q}+\frac{\beta_{q}}{2}$, while $s^{zp}_{i}(v_{i})=p_{i}^{zp}(v_{i})+jq_{i}^{zp}(v_{i})$. The new coefficients in the ZP model are obtained by the binomial approximation method. The ZP approximation is valid as long as the voltage is close enough to the nominal value i.e., while $|v_{i}-v_{nom}|\leq 0.1$, the ZP approximation is valid [31] ($v_{nom}$ is the nominal voltage).

### v6

@@ Abstract
The widespread integration of IoT-controllable devices (e.g., smart EV charging stations and heat pumps) into modern power systems enhances capabilities but introduces critical cybersecurity risks. Specifically, these devices are susceptible to load-altering attacks (LAAs) that can compromise power system safety. This paper quantifies the impact of LAAs on nodal voltage constraint violations in distribution networks (DNs). We first present closed-form expressions to analytically characterize LAA effects and quantify the minimum number of compromised devices for a successful LAA. Based on these insights, we propose a reactive defense mechanism that mitigates LAAs through DN reconfiguration. To address strategic adversaries, we then formulate defense strategies using a non-cooperative sequential game, which models the knowledgeable and strategic attacker, accounting for the worst-case scenario and enabling the reactive defender to devise an efficient and robust defense. Further, our formulation also accounts for uncertainties in attack localization. A novel Bayesian optimization approach is introduced to compute the Stackelberg equilibrium, significantly reducing computational burden efficiently. The game-theoretic strategy effectively mitigates the attack’s impact while ensuring minimal system reconfiguration.
Nomenclature
  • Parameters
  • $\alpha^{\prime}_{p},\gamma^{\prime}_{p}$
    Active ZP load coefficients
  • $\alpha^{\prime}_{q},\gamma^{\prime}_{q}$
    Reactive ZP load coefficients
  • $\alpha_{p},\beta_{p},\gamma_{p}$
    Active ZIP load coefficients
  • $\alpha_{q},\beta_{q},\gamma_{q}$
    Reactive ZIP load coefficients
  • $\overline{s_{\text{der},i}}$
    Maximum output capacity of DER in bus i
  • I
    Identity matrix
  • $b_{ij}^{pre}$
    Component of the adjacency matrix before reconfiguration in $i^{th}$ row and $j^{th}$ column
  • $M$
    Disjunctive parameter
  • NN
    Number of buses of the distribution system
  • $p_{d},q_{d}$
    Nominal active and reactive power demands by a single attacked device
  • $p_{i}^{l},q_{i}^{l}$
    Active and reactive load demands in bus $i$
  • $r_{ij},x_{ij}$
    Resistance and reactance of the line from bus $i$ to $j$
  • $v_{nom}$
    Nominal voltage
  • Sets
  • $\mathcal{D}_{i}$
    Set of buses engaged in the unique path connecting the bus $i$ to the root bus
  • $\mathcal{L}\setminus\mathcal{L}^{s}$
    Set of lines without switches
  • $\mathcal{L}^{s}$
    Set of lines with switches
  • $\mathcal{N}\setminus\mathcal{N}^{f}$
    Subset of buses which are not substations
  • $\mathcal{N}^{a}$
    Subset of buses under attack
  • $\mathcal{N}^{f}$
    Subset of buses which are substation
  • $\mathcal{N}^{\text{DER}}$
    Set of buses with DER
  • Variables
  • $\hat{v}_{ji}$
    Auxiliary voltage variable for MILP
  • $\Omega,\Omega^{\prime},\Omega^{\prime\prime}$
    Coefficients matrices for the proposed closed-form expressions
  • $\pi_{i}$
    Parent bus of the bus $i$
  • $\sigma_{\ell}$
    Probability of the bus $\ell$ to be under attack
  • U
    Vector of the square of voltages of the system while no LAA
  • $\textbf{U}^{A}$
    Vector of square of voltages of the system under LAA
  • $a^{*},r(a^{*})$
    Players actions in equilibrium
  • $B$
    Adjacency matrix of the distribution network after reconfiguration
  • $b_{ij}$
    Component of $B$ in $i^{th}$ row and $j^{th}$ column
  • $c(n)$
    Number of attacked devices for a “critical attack” in bus $n$
  • $I_{\pi_{i},i}$
    Current flowing through the branch $(\pi_{i},i)\in\mathcal{L}$
  • $n_{att}$
    The bus under attack
  • $p^{a},q^{a}$
    Active and reactive powers raised by LAA
  • $p_{i}^{0}$, $p_{i}^{0}$
    Active and reactive load demands at rated voltage in bus $i$
  • $p_{i}^{f},q_{i}^{f}$
    Active and reactive powers flowing from substation bus $i$
  • $p_{\text{der},i},q_{\text{der},i}$
    Active and reactive power output of DER in bus i
  • $p_{ij},q_{ij}$
    Active and reactive power flows from bus $i$ to $j$
  • $r(a)$
    Follower’s best response to attacker’s strategy
  • $s_{i}$
    Apparent power in bus $i$
  • $s_{i}^{\text{zip}},s_{i}^{\text{zp}}$
    Apparent power in bus $i$ with ZIP and ZP load models
  • $v_{i}$
    Voltage in bus $i$
@@ I Introduction
Internet-of-Things (IoT) devices offer enhanced end-user convenience, improved efficiency and flexibility to power systems for load peak management, which has driven a notable surge in their adoption. However, beyond their evident benefits, these devices also present potential vulnerabilities, serving as entry points for cyber attackers to compromise the security of power systems. Specifically, load-altering attacks (LAA) in power networks with high IoT-enabled device penetration pose a significant cybersecurity threat [2, 3, 4, 5].
The concept of LAAs was first introduced in [2]. In this work, adversaries turn a group of IoT-controllable electrical loads into bots, which are then turned on and off simultaneously to harm the stability of the system. The manipulation of loads disrupts the balance between power generation and demand, leading to frequency instabilities in transmission networks [6, 4]. In distribution networks (DNs), LAA can result in elevated line flows, causing higher voltage drops, leading to voltage constraint violation [7]. Furthermore, the data required to launch a successful LAA can be obtained from publicly available information [8] or bypassed [9].
@@ I-A Literature Survey
Attack Impact Analysis and Viability: Researchers in [4] investigated the impact of LAA on transmission systems and identified several effects, including line failure, unsafe frequency deviation, disruption in grid restarting, and increased operational cost. The research in [10] examined the effects of LAA under a more realistic setting. This setting included protection schemes (such as N1N-1 security) and load-shedding schemes. The study showed that even in this context, LAA can still cause outages and islands. Additionally, a recent work [11] expands the LAA model and launches the attack on high-wattage IoT devices and distributed energy resources (DERs) simultaneously. One notable result from their work is that attacking scattered devices causes a lower impact than attacks on devices in a concentrated area.
Reference [3] proposed the dynamic LAA (DLAA), in which the adversary continuously toggles the compromised load devices on and off, guided by a feedback control loop in response to the system’s frequency fluctuations. In [5], an analytical framework was introduced to analyze the impact of LAA on transmission systems and identify the nodes from which an attacker can launch the most effective attacks using the theory of second-order dynamical systems. In [12], a rare-event sampling algorithm was proposed that uncovers the spatial and temporal distribution of impactful DLAAs while considering the security constraint of N1N-1.
The growing popularity of electric vehicles (EVs) and the spread of EV charging stations (EVCS) make them a potential target for the LAA. Reference [13] explored the weak points of EVCS, such as firmware flaws, vulnerabilities in management systems, and the security scarcity of mobile apps that allow adversaries to launch LAAs. Subsequently, to detect such attacks, they propose model-based and data-driven approaches.
Attack Mitigation: Another stream of research investigates the mitigation of LAAs. The existing mitigation methods can be categorized into: i) offline, or ii) online methods.
@@
Offline methods: Offline defenses try to install preventive measures to stop the impact of LAA. For instance, [14] proposed algorithms to determine the operating points for generators in a way to prevent line overloads caused by potential botnet-type attacks against IoT devices. Reference [3] presented a mitigation framework based on securing the most critical loads. This method found the minimum magnitude of loads needed to be protected in order to guarantee frequency stability in the event of DLAAs. [19] proposed a zero-sum Stackelberg game formulation to install reactive power compensation (RPC) to reduce the impact of the attack. While the works above focus on transmission systems, [7] introduced a mitigation approach tailored specifically for DNs. Their research focuses on identifying optimal locations for deploying soft open points (SOPs) and refining their operation to mitigate the effects of attacks on voltage deviations.
@@
Online methods: Despite the effectiveness of the offline methods, these measures may be too costly as the preventive features must be enabled irrespective of whether an attack occurs or not (e.g., uneconomic generator operating points to cover for LAAs). Online methods, on the other hand, involve determining defensive actions to counter the effects of LAAs once the attack is launched, via reactive measures. In [15], a cyber-resilient economic dispatch method is introduced to mitigate LAAs based on altering the frequency droop control parameters of inverter-based resources to counter the destabilizing effects of LAAs. [16] proposed a framework in which electric vehicles are designed as feedback controllers that can mitigate the impact of LAA based on H2H-2 and $H$-$\infty$ norms. To analyze the manoeuvres of a strategic attacker initiating DLAA, [17] introduced a multi-stage game approach. In this game, the defensive actions involve load shedding, and the ultimate objective is to achieve a strategic balance between DLAA and the necessary amount of load shedding, reaching a Nash equilibrium (NE). [18] formulated the cybersecurity of the power system in both cyber and physical layers as a game-theoretic formulation.
Existing research on LAAs has largely focused on transmission grids. However, cybersecurity research on DNs extends beyond LAAs. One group of studies investigates attack models and vulnerabilities in DNs, aiming to identify potential threats and prepare for them as the pre-attack stage. These include assessing load redistribution attacks in unbalanced DNs [20], proposing a dynamic false data injection attack [21], and highlighting critical operations and components that are prone to cyber-physical attacks [22].
A second group of studies addresses the during-attack stage by developing detection and mitigation frameworks. This includes hierarchical methods leveraging waveform measurements [23], unsupervised adversarial autoencoders for attack detection [24], and vehicle-to-grid (V2G) voltage control schemes for attack mitigation [25].
A third group focuses on the post-attack stage. For example, [26] proposes a high-level three-step grid restoration framework consisting of (1) post-attack detection and localization, (2) isolation and pre-recovery, and (3) recovery and assessment. Furthermore, [27] studies optimal crew routing during the cyber-recovery phase.
Distinct from these works, this paper first analyzes the impact of LAAs on DNs. Based on the insights gained and through a game-theoretic study, a bespoke defense mechanism is proposed that leverages the existing infrastructure of the grid.
Differences of Attack on Transmission and Distribution Grids: Due to the radial structure of DNs, an attacker can cause severe voltage drops by compromising a small number of devices at deep (leaf) nodes. For example, in [10], the authors conclude that an effective LAA on the PowerWorld 9-bus transmission grid needs to alter 30% of its total load, which is equivalent to 24682 MW (the power consumed by millions of air conditioners). However, the analysis in this paper shows that a successful attack on the 33-bus DN can be as small as compromising less than 100 air conditioners. This spatial sensitivity does not exist in the same form in transmission systems. Additionally, this spatial sensitivity enables us to propose a reconfiguration-based [28] defense in the game-theoretic analysis, whereas previous transmission-level games focus on load shedding or generator dispatch. These diverse defense models also stem from the different threatened stabilities. In transmission grids, LAA primarily threatens frequency stability, often necessitating frequency-based or generator-side mitigation strategies. However, in DNs, LAA mainly affects voltage regulation.
Furthermore, novel DNs have higher dependence on the information technology section, and this makes them a prime target for attackers [22]. Transmission grids typically have robust protection infrastructures; however, a single successful breach can have catastrophic consequences, such as widespread blackouts. In contrast, cyberattacks on DNs can be launched more frequently and with greater feasibility, though a single incident is less likely to cause large-scale disruption.
@@ I-B Research Gap and Contributions
Despite the growing threat of LAAs, existing research mainly focuses on transmission systems, with limited works addressing their impact or mitigation in DNs. Moreover, few studies account for strategic attackers who adapt based on system defenses. This paper addresses these gaps by analyzing the effects of LAA in DNs and proposing a game-theoretic defense strategy based on network reconfiguration.
This paper significantly extends our preliminary work [1] and addresses these challenges and proposes a defense strategy specifically tailored for DN, highlighting several conceptual innovations:
$\bullet$ Deriving closed-form expressions to assess how LAA affects nodal voltages and to quantify the minimum number of IoT-controlled devices needed to cause voltage violations in DNs with voltage-dependent loads.
$\bullet$ Developing a Stackelberg game to model interactions between a strategic attacker and the DN operator, incorporating uncertainty in attack localization.
$\bullet$ Introducing a Bayesian optimization method to compute the Stackelberg equilibrium efficiently, significantly reducing the computational burden of finding optimal reconfiguration strategies.
These contributions collectively enable a practical and scalable framework for analyzing and mitigating LAA in DNs.
@@ II-A DN Model
The DN under study is represented by the connected directed graph $G=\{\mathcal{N},\mathcal{L}\}$, where $\mathcal{N}=\{1,2,\ldots,N\}$ denotes the set of buses and $\mathcal{L}$ denotes the set of branches. This graph has a radial structure, hence it is a tree. Except for bus $1$, which is the root, each bus is referred to as the ‘child’ of its parent bus, which is the adjacent bus closer to bus $1$ by one branch. Thus, the set of branches is defined as $\mathcal{L}=\{(\pi_{i},i)\ |\ \pi_{i},\ i\in\mathcal{N}\}$. In this configuration, bus $1$ represents the generator bus. We denote by $\mathcal{D}_{k}$ the set of buses which form the unique path connecting bus $1$ to bus $k$, excluding bus $1$ and including bus $k$. The depth of each bus represents the distance in terms of the number of branches between that bus and the root bus.
@@ II-B1 ZIP Load Model
ZIP and exponential load models are the most implemented ones in the industry. Experimental values for the voltage dependency of loads are captured and fitted in the ZIP model in [29], which we integrate into our formulations. The power demand under the ZIP load model is given in [30] as follows:
$s^{\text{zip}}_{i}(v_{i})=p^{l}_{i}(\alpha_{p}+\beta_{p}v_{i}+\gamma_{p}v^{2}_{i})+jq^{l}_{i}(\alpha_{q}+\beta_{q}v_{i}+\gamma_{q}v^{2}_{i}),$ (1)
where $\alpha_{k}+\beta_{k}+\gamma_{k}=1$, where $k=\{p,q\}.$ The ZIP load model captures the voltage dependency of real-world loads.
@@ II-B2 ZP Approximation
Based on (1), the ZIP model is a function of both $v_{i}$ and $v^{2}_{i}$. This causes the optimization tasks involving the power flow in the presence of ZIP loads to become nonconvex and complex. To tackle this problem, [31] provided an approximate model for ZIP loads given by
$s^{\text{zp}}_{i}(v_{i})=p^{l}_{i}(\alpha^{\prime}_{p}+\gamma^{\prime}_{p}v_{i}^{2})+jq^{l}_{i}(\alpha^{\prime}_{q}+\gamma^{\prime}_{q}v^{2}_{i}),$ (2)
where $\alpha^{\prime}_{p}=\alpha_{p}+\frac{\beta_{p}}{2}$ $\alpha^{\prime}_{q}=\alpha_{q}+\frac{\beta_{q}}{2}$ $\gamma^{\prime}_{p}=\gamma_{p}+\frac{\beta_{p}}{2}$ , and $\alpha^{\prime}_{q}=\gamma_{q}+\frac{\beta_{q}}{2}$, while $s^{\text{zp}}_{i}(v_{i})=p_{i}^{\text{zp}}(v_{i})+jq_{i}^{\text{zp}}(v_{i})$. The new coefficients in the ZP model are obtained by the binomial approximation method. The ZP approximation is valid as long as the voltage is close enough to the nominal value, i.e., while $|v_{i}-v_{nom}|\leq 0.1$, the ZP approximation is valid [31] ($v_{nom}$ is the nominal voltage). To further clarify this, we illustrate the dependence of apparent power on voltage using the ZIP and ZP models for an air conditioner in Figure 2. As shown, when the voltage is close to 1 p.u., the ZP model produces values very similar to those of the ZIP model. According to [31], the ZP approximation, regardless of the modeled device, is sufficiently accurate when the voltage lies between 0.9 p.u. and 1.1 p.u. (the yellow zone). In our work, according to constraint (8), which is used for all optimizations, the voltage is constrained in the range [0.95, 1.05] p.u. (highlighted in green in the figure), which implies that the ZP approximation is valid for our setting.
Figure 2: Power demand dependence of an air conditioner on voltage in accordance with ZIP and ZP models. The green zone illustrates the voltage constraints in this research, and the yellow zone is the accurate region of the ZP approximation